OpenStudy (anonymous):

multiply x 1/4 times x 5/8

OpenStudy (anonymous):

Are you asking on how to solve it, or how to write it in words?

OpenStudy (anonymous):

how to solve it

OpenStudy (anonymous):

\[x \frac{ 1 }{ 4 } * x \frac{ 5 }{ 8 }\]

OpenStudy (anonymous):

I think u meant1/4x times 5/8x not x1/4 time x5/8

OpenStudy (anonymous):

its actually x^1/4 times x^5/8

OpenStudy (anonymous):

Solve that equation I gave you, and tell me what you get.

OpenStudy (anonymous):

x5/40

OpenStudy (anonymous):

x times x is x^2

OpenStudy (anonymous):

not quite. But you're right on the answer will be a fraction. \[5*1=5\] \[x*x=x^2\] \[8*4=32\] Therefore, \[\frac{ 5x^2 }{ 32 }\]

OpenStudy (anonymous):

Thank you @lacrosseplayer22 for clearing that up.

OpenStudy (anonymous):

np:)

OpenStudy (anonymous):

thank you guys but the closest answer choice i have is x^5/32

OpenStudy (anonymous):

That'll do.

OpenStudy (anonymous):

You don't have a square root answer?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

thank you